# Adaptive Variance Reduction for Stochastic Optimization under Weaker Assumptions

### v1

@@ 1 Introduction
We give an affirmative answer to the above question by devising a novel optimal Adaptive STORM method (Ada-STORM). The learning rate of our algorithm is set to be inversely proportional to a specific power of the iteration number $T$ in the initial iterations, and then changes adaptively to the cumulative sum of past gradient estimations. In this way, we are able to adjust the learning rate dynamically according to the property of stochastic gradients, and ensure a small learning rate in the beginning. Leveraging this strategy, Ada-STORM achieves an optimal convergence rate of $\mathcal{O}(T^{-1/3})$ for non-convex functions. Notably, our analysis does not require the function to have bounded values and bounded gradients, which is a significant advancement over existing methods Levy et al. (2021); Liu et al. (2022). Additionally, our convergence rates do not contain the extra $\mathcal{O}(\log T)$ term, which is often present in STORM-based methods Cutkosky and Orabona (2019); Liu et al. (2022). To highlight the versatility of our approach and its potential impact in the field of stochastic optimization, we further extend our technique to develop optimal adaptive methods for compositional optimization.
@@ 3.2 The proposed method
The core idea of STORM lies in a carefully devised variance reduced estimator $\mathbf{v}_{t}$, which effectively tracks the gradient $\nabla f(\mathbf{x}_{t})$. For the first iteration ($t=1$), we set $\mathbf{v}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}\nabla f(\mathbf{x}_{1};\xi_{1}^{i})$, which is estimated within a batch $B_{0}=T^{1/3}$. Note that we use large batch only in the first iteration, and constant batch size in other iterations. In subsequent iterations ($t>2$), estimator $\mathbf{v}_{t}$ is defined as:
@@
Remark: To ensure that $\mathbb{E}[\left\|\nabla f(\mathbf{x}_{\tau})\right\|]\leq\epsilon$, the overall complexity is $\mathcal{O}(\epsilon^{-3})$, which is known to be optimal up to constant factors Arjevani et al. (2019). Compared with existing STORM-based algorithms Cutkosky and Orabona (2019); Levy et al. (2021); Liu et al. (2022), our method does not have the extra $\mathcal{O}(\log T)$ term in the convergence rate, and our analysis does not require bounded gradients or bounded function values assumptions. Also note that the selection of $\alpha$ does not affect the order of $T$, and larger $\alpha$ leads to better dependence on parameter $L$. Considering we require that $30<\alpha<1/$ 3, we can simply set $\alpha=0.3$ in practice.
@@ 3.3 The doubling trick
While we have attained the optimal convergence rate using the proposed adaptive STORM method, it requires knowing the total number of iterations $T$ in advance. Here, we show that we can avoid this requirement by using the doubling trick, which divides the algorithm into several stages and increases the iteration number in each stage gradually. Specifically, we design a multi-stage algorithm over $k=\{1,2,\cdots,K\}$ stages. In each stage $k$, the STORM algorithm is executed for $2^{k-1}$ iterations, effectively doubling the iteration numbers after each stage. In any step $t$, we first identify the current stage as $1+\lfloor\log t\rfloor$ and then calculate the iteration number for this stage as $I_{t}=2^{\lfloor\log t\rfloor}$. Then, we can set the hyper-parameters as:
@@ 4 Extension to stochastic compositional optimization
To demonstrate the broad applicability of our proposed technique, we extend it to stochastic compositional optimization Wang et al. (2017a, b); Yuan et al. (2019); Zhang and Xiao (2019, 2021); Jiang et al. (2023), formulated as:
@@
After that, we apply gradient descent using the gradient estimator $\mathbf{v}_{t}$. The whole algorithm is presented in Algorithm 2, and hyper-parameters are set the same as in equation (3). For the first iteration, we simply set $\mathbf{u}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}g(\mathbf{x}_{1};\zeta_{1}^{i})$ and $\mathbf{v}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}\nabla f(\mathbf{u}_{1};\xi_{1}^{i})\nabla g(\mathbf{x}_{1};\zeta_{1})$, where $B_{0}=T^{1/3}$. Next, we list common assumptions used in the literature of compositional optimization Wang et al. (2017a, b); Yuan et al. (2019); Zhang and Xiao (2019, 2021); Jiang et al. (2023).
@@ 5 Adaptive variance reduction for finite-sum optimization
Algorithm 3 STORM for Finite-sum Optimization
@@
Remark: Our result matches the lower bound for non-convex finite-sum problems Fang et al. (2018); Li et al. (2021), and makes an improvement over the existing adaptive method Kavis et al. (2022) by eliminating the $\mathcal{O}(\log(nT))$ factor.
Remark: We can avoid storing past gradients by following the SVRG method (Johnson and Zhang, 2013; Zhang et al., 2013) to compute the full gradient periodically and incorporate it into STORM estimator. This strategy maintains the same optimal rate and more details are deferred to Appendix D.1.
@@ 7 Conclusion
In this paper, we propose an adaptive STORM method to achieve the optimal convergence rate for non-convex functions. Compared with existing methods, our algorithm requires weaker assumptions and does not have the additional $\mathcal{O}(\log T)$ term in the convergence rate. The proposed technique can also be employed to develop optimal adaptive algorithms for compositional optimization. Furthermore, we investigate an adaptive method for non-convex finite-sum optimization, obtaining an improved convergence rate of $\mathcal{O}(n^{1/4}T^{-1/2})$. Given that STORM algorithm has already been used in many areas such as bi-level optimization Yang et al. (2021), federated learning Das et al. (2022), min-max optimization Xian et al. (2021), finite-sum compositional optimization Jiang et al. (2022), etc., the proposed methods may also inspire the development of adaptive algorithms in these fields.
@@ Appendix A Proof of Theorem 1
which indicate that
@@
with $C_{0}$ defined in equation (14). We find that larger $\alpha$ leads to better dependence on $L$. For $\alpha\to\frac{1}{3}$, we can obtain that
@@ Appendix C Proof of Theorem 3
$\leq(1-\beta)\left\|\mathbf{v}_{t}-G_{t}\right\|^{2}+2\beta^{2}\mathbb{E}_{\xi_{t+1},\zeta_{t+1}}\left[\left\|\nabla f(\mathbf{u}_{t+1};\xi_{t+1})\nabla g(\mathbf{x}_{t+1};\zeta_{t+1})-\nabla f(\mathbf{u}_{t+1})\nabla g(\mathbf{x}_{t+1})\right\|^{2}\right]$
@@
$\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}\left\|\nabla F(\mathbf{x}_{t})-\mathbf{v}_{t}\right\|^{2}\right]$
$\leq$ $26C^{2}L^{2}\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}\left\|\mathbf{u}_{t}-g(\mathbf{x}_{t})\right\|^{2}\right]+\frac{8C^{2}L^{2}+24C^{4}L^{2}}{\beta}\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}^{3}\left\|\mathbf{v}_{t}\right\|^{2}\right]+(2+8C^{2}+24C^{2}L^{2})\sigma^{2}.$
@@
$\leq$ $26C^{2}L^{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\mathbf{u}_{t}-g(\mathbf{x}_{t})\right\|^{2}\right]+\frac{8C^{2}L^{2}+24C^{4}L^{2}}{\beta}\mathbb{E}\left[\sum_{t=1}^{T}\eta_{t}^{2}\left\|\mathbf{v}_{t}\right\|^{2}\right]+(2+8C^{2}+24C^{2}L^{2})\sigma^{2}T^{1/3}$
@@ Appendix D Proof of Theorem 4
$\begin{split}&\quad\ \mathbb{E}_{i_{t+1}}\left[\left\|\nabla F(\mathbf{x}_{t+1})-\mathbf{v}_{t+1}\right\|^{2}\right]\\ &=\mathbb{E}_{i_{t+1}}\left[\left\|(1-\beta)\mathbf{v}_{t}+\beta\mathbf{z}_{t+1}+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right)-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &=\mathbb{E}_{i_{t+1}}\left[\|(1-\beta)(\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t}))+\beta(\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1}))\right.\\ &\quad\left.+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right)\|^{2}\right]\\ &\leq\left\|(1-\beta)(\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t}))\right\|^{2}\\ &\quad+\mathbb{E}_{i_{t+1}}\left[\left\|\beta(\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1}))+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right)\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\quad+2(1-\beta)^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\quad+2(1-\beta)^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]+2L^{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}\\ &\leq(1-\beta)\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]+2L^{2}\eta_{t}^{2}\left\|\mathbf{v}_{t}\right\|^{2}.\end{split}$ (17)
@@
$\leq$ $\mathbb{E}_{i_{t+1}}\left[(1+2n)L^{2}\eta_{t}^{3}\left\|\mathbf{v}_{t}\right\|^{2}+\eta_{t}(1+\frac{1}{2n})\left((1-\frac{1}{n})\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t})-g_{t}^{i_{t+1}}\right\|^{2}+\frac{1}{n}\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right\|^{2}\right]\right)$
@@
$\sum_{t=1}^{T}\eta_{t}^{2}L\left\|\mathbf{v}_{t}\right\|^{2}=$ ${L^{2}}\sum_{t=1}^{T}\frac{\left\|\mathbf{v}_{t}\right\|^{2}}{n^{1-\alpha}\left(\sum_{i=1}^{t}\left\|\mathbf{v}_{i}\right\|^{2}\right)^{2\alpha}}$
@@
$C_{3}=2\Delta_{F}+\frac{14\alpha}{1-\alpha}\left(\frac{2L^{2}}{1-3\alpha}\left(\frac{24-72\alpha}{1-\alpha}\right)^{\frac{1-3\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{2\alpha}}+\frac{\alpha}{1-\alpha}\left(\frac{L}{1-2\alpha}\left(\frac{4-8\alpha}{1-\alpha}\right)^{\frac{1-2\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{\alpha}}.$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\left\|\nabla F(\mathbf{x}_{t})-\mathbf{v}_{t}\right\|^{2}\right]\leq\frac{\alpha\sqrt{n}}{1-\alpha}\left(\frac{14L^{2}}{1-2\alpha}\left(\frac{2-4\alpha}{(1-\alpha)L}\right)^{\frac{1-2\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{\alpha}}+\frac{n^{\frac{\alpha}{2}}L}{2}\mathbb{E}\left[\left(\sum_{t=1}^{T}\left\|\mathbf{v}_{t}\right\|^{2}\right)^{1-\alpha}\right].$

### v2

@@ 1 Introduction
We give an affirmative answer to the above question by devising a novel optimal Adaptive STORM method (Ada-STORM). The learning rate of our algorithm is set to be inversely proportional to a specific power of the iteration number $T$ in the initial iterations, and then changes adaptively based on the cumulative sum of past gradient estimations. In this way, we are able to adjust the learning rate dynamically according to the property of stochastic gradients, and ensure a small learning rate in the beginning. Leveraging this strategy, Ada-STORM achieves an optimal convergence rate of $\mathcal{O}(T^{-1/3})$ for non-convex functions. Notably, our analysis does not require the function to have bounded values and bounded gradients, which is a significant advancement over existing methods Levy et al. (2021); Liu et al. (2022). Additionally, our convergence rate does not contain the extra $\mathcal{O}(\log T)$ term, which is often present in STORM-based methods Cutkosky and Orabona (2019); Liu et al. (2022). To highlight the versatility of our approach and its potential impact in the field of stochastic optimization, we further extend our technique to develop optimal adaptive methods for compositional optimization.
@@ 3.2 The proposed method
The core idea of STORM lies in a carefully devised variance reduced estimator $\mathbf{v}_{t}$, which effectively tracks the gradient $\nabla f(\mathbf{x}_{t})$. For the first iteration ($t=1$), we set $\mathbf{v}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}\nabla f(\mathbf{x}_{1};\xi_{1}^{i})$, which is estimated within a batch $B_{0}=T^{1/3}$. Note that we use large batch only in the first iteration, and constant batch size in other iterations. In subsequent iterations ($t\geq 2$), estimator $\mathbf{v}_{t}$ is defined as:
@@
Remark: To ensure that $\mathbb{E}[\left\|\nabla f(\mathbf{x}_{\tau})\right\|]\leq\epsilon$, the overall complexity is $\mathcal{O}(\epsilon^{-3})$, which is known to be optimal up to constant factors Arjevani et al. (2019). Compared with existing STORM-based algorithms Cutkosky and Orabona (2019); Levy et al. (2021); Liu et al. (2022), our method does not have the extra $\mathcal{O}(\log T)$ term in the convergence rate, and our analysis does not require bounded gradients or bounded function values assumptions. Also note that the selection of $\alpha$ does not affect the order of $T$, and larger $\alpha$ leads to better dependence on parameter $L$ and worse reliance on parameters $\Delta$ and $\sigma$. Considering we require that $30<\alpha<1/$ 3, we can simply set $\alpha=0.3$ in practice.
@@ 3.3 The doubling trick
While we have attained the optimal convergence rate using the proposed adaptive STORM method, it requires knowing the total number of iterations $T$ in advance. Here, we show that we can avoid this requirement by using the doubling trick, which divides the algorithm into several stages and increases the iteration number in each stage gradually. Specifically, we design a multi-stage algorithm over $k=\{1,2,\cdots,K\}$ stages. At the beginning of each new stage, we reset $\mathbf{x}_{t}=\mathbf{x}_{0}$. In each stage $k$, the STORM algorithm is executed for $2^{k-1}$ iterations, effectively doubling the iteration numbers after each stage. In any step $t$, we first identify the current stage as $1+\lfloor\log t\rfloor$ and then calculate the iteration number for this stage as $I_{t}=2^{\lfloor\log t\rfloor}$. Then, we can set the hyper-parameters as:
@@ 4 Extension to stochastic compositional optimization
To demonstrate the broad applicability of our proposed technique, we extend it to stochastic compositional optimization Wang et al. (2017a, b); Yuan et al. (2019); Zhang and Xiao (2019, 2021); Jiang et al. (2023, 2024a), formulated as:
@@
After that, we apply gradient descent using the gradient estimator $\mathbf{v}_{t}$. The whole algorithm is presented in Algorithm 2, and hyper-parameters are set the same as in equation (3). For the first iteration, we simply set $\mathbf{u}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}g(\mathbf{x}_{1};\zeta_{1}^{i})$ and $\mathbf{v}_{1}=\sum_{i=1}^{B_{0}}\frac{1}{B_{0}}\nabla f(\mathbf{u}_{1};\xi_{1}^{i})\nabla g(\mathbf{x}_{1};\zeta_{1})$, where $B_{0}=T^{1/3}$. Next, we list common assumptions used in the literature of compositional optimization Wang et al. (2017a, b); Yuan et al. (2019); Zhang and Xiao (2019, 2021).
@@
Remark: In Assumption 4, we further require standard Lipschitz continuity assumption, which is essential and widely required in the literature for stochastic compositional optimization Wang et al. (2017b); Yuan et al. (2019); Jiang et al. (2022a, b). This assumption is inherently introduced by the compositional optimization itself rather than by our adaptive techniques.
@@ 5 Adaptive variance reduction for finite-sum optimization
Algorithm 3 STORM for Finite-sum Optimization (SAG-type)
@@
Remark: Our result matches the lower bound for non-convex finite-sum problems Fang et al. (2018); Li et al. (2021), and makes an improvement over the existing adaptive method, i.e., AdaSpider Kavis et al. (2022). Specifically, the convergence rate of the AdaSpider algorithm is $\mathcal{O}\left(n^{1/4}T^{-1/2}\left(L^{2}+\Delta_{F}\right)\cdot\log\left(1+nTL\right)\right)$, and our result is better than theirs when $\frac{1}{4}<\alpha<\frac{1}{3}$.
We can avoid storing past gradients by following the SVRG method (Johnson and Zhang, 2013; Zhang et al., 2013) to compute the full gradient periodically and incorporate it into STORM estimator. Instead of storing the past gradients as in SAG algorithm, we can avoid this storage cost by incorporating elements from the SVRG method. Specifically, we compute a full batch gradient at the first step and every $I$ iteration (we set $I=n$):
$\nabla f(\mathbf{x}_{\tau})=\frac{1}{n}\sum_{i=1}^{n}\nabla f_{i}(\mathbf{x}_{\tau}).$
Note that the first three terms match the original STORM estimator, and the last term, inspired from SVRG, deals with the finite-sum structure. Compared with equation (8) in Algorithm 3, the difference is that we use $\left(\nabla f_{i_{t}}(\mathbf{x}_{\tau})-\nabla f(\mathbf{x}_{\tau})\right)$ instead of $\left(g_{t}^{i_{t}}-\frac{1}{n}\sum_{i=1}^{n}g_{t}^{i}\right)$ in the last term. The detailed procedure is outlined in Algorithm 4. This strategy maintains the same optimal rate, as stated below:
Theorem 5
Under Assumptions 7 and 8, our Algorithm 4 guarantees that:
Remark: The obtained convergence rate is in the same order as the results in Theorem 4, and Algorithm 4 does not require storing past gradients anymore.
Algorithm 4 STORM for Finite-sum Optimization (SVRG-type)
1:  Input: time step $T$, initial point $\mathbf{x}_{1}$
2:  for time step $t=1$ to $T$ do
3:     if $t\mod I==0$ then
5:     end if
6:     Sample $i_{t}$ randomly from $\{1,2,\cdots,n\}$
7:     Compute $\mathbf{v}_{t}$ according to equation (10)
9:  end for
10:  Select $\tau$ uniformly at random from $\{1,\ldots,T\}$
11:  Return $\mathbf{x}_{\tau}$
@@ 7 Conclusion
In this paper, we propose an adaptive STORM method to achieve the optimal convergence rate for non-convex functions. Compared with existing methods, our algorithm requires weaker assumptions and does not have the additional $\mathcal{O}(\log T)$ term in the convergence rate. The proposed technique can also be employed to develop optimal adaptive algorithms for compositional optimization. Furthermore, we investigate an adaptive method for non-convex finite-sum optimization, obtaining an improved convergence rate of $\mathcal{O}(n^{1/4}T^{-1/2})$. Given that STORM algorithm has already been used in many areas such as bi-level optimization Yang et al. (2021), federated learning Das et al. (2022), min-max optimization Xian et al. (2021), sign-based optimization Jiang et al. (2024b), etc., the proposed methods may also inspire the development of adaptive algorithms in these fields.
Acknowledgements
This work was partially supported by National Key R&D Program of China (2021ZD0112802), NSFC (62122037), and the Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. KYCX24_0231).
@@ Appendix A Proof of Theorem 1
$\leq\mathbb{E}_{\xi_{t+1}}\left[(1-\beta)^{2}||\mathbf{v}_{t}-\nabla f(\mathbf{x}_{t})||^{2}\right]+\mathbb{E}_{\xi_{t+1}}\left[||\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{x}_{t+1})\right.$
$\qquad\left.+\nabla f(\mathbf{x}_{t+1};\xi_{t+1})-\nabla f(\mathbf{x}_{t};\xi_{t+1})+\beta\left(\nabla f(\mathbf{x}_{t};\xi_{t+1})-\nabla f(\mathbf{x}_{t})\right)||^{2}\right]$
$\leq\mathbb{E}_{\xi_{t+1}}\left[(1-\beta)^{2}||\mathbf{v}_{t}-\nabla f(\mathbf{x}_{t})||^{2}\right]$
$\quad+2\mathbb{E}_{\xi_{t+1}}\left[||\nabla f(\mathbf{x}_{t})-\nabla f(\mathbf{x}_{t+1})+\nabla f(\mathbf{x}_{t+1};\xi_{t+1})-\nabla f(\mathbf{x}_{t};\xi_{t+1})||^{2}\right]$
$\quad+2\mathbb{E}_{\xi_{t+1}}\left[||\beta\left(\nabla f(\mathbf{x}_{t};\xi_{t+1})-\nabla f(\mathbf{x}_{t})\right)||^{2}\right]$
@@
which indicates the following by applying Jensen’s inequality:
@@
Also, because of Jensen’s inequality, we have:
$\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{s-1}||\mathbf{v}_{t}||\right]$ $\leq\sqrt{\mathbb{E}\left[\left(\frac{1}{T}\sum_{t=1}^{s-1}||\mathbf{v}_{t}||\right)^{2}\right]}=\sqrt{\mathbb{E}\left[\frac{1}{T^{2}}\left(\sum_{t=1}^{s-1}||\mathbf{v}_{t}||\right)^{2}\right]}$
$\leq\sqrt{\mathbb{E}\left[\frac{s}{T^{2}}\sum_{t=1}^{s-1}||\mathbf{v}_{t}||^{2}\right]}\leq\sqrt{\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{s-1}||\mathbf{v}_{t}||^{2}\right]}$
$\leq\sqrt{\frac{1}{T}C_{0}T^{1/3}}=\sqrt{C_{0}}T^{-1/3}.$
@@
with $C_{0}$ defined in equation (15). We find that larger $\alpha$ leads to better dependence on $L$ and worse reliance on parameters $\Delta$ and $\sigma$. For $\alpha\to\frac{1}{3}$, we can obtain that
@@ Appendix C Proof of Theorem 3
$\leq(1-\beta)\left\|\mathbf{v}_{t}-G_{t}\right\|^{2}$
$\qquad+2\beta^{2}\mathbb{E}_{\xi_{t+1},\zeta_{t+1}}\left[\left\|\nabla f(\mathbf{u}_{t+1};\xi_{t+1})\nabla g(\mathbf{x}_{t+1};\zeta_{t+1})-\nabla f(\mathbf{u}_{t+1})\nabla g(\mathbf{x}_{t+1})\right\|^{2}\right]$
@@
$\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}\left\|\nabla F(\mathbf{x}_{t})-\mathbf{v}_{t}\right\|^{2}\right]\leq$ $26C^{2}L^{2}\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}\left\|\mathbf{u}_{t}-g(\mathbf{x}_{t})\right\|^{2}\right]$
$+$ $\frac{8C^{2}L^{2}+24C^{4}L^{2}}{\beta}\sum_{t=1}^{T}\mathbb{E}\left[\eta_{t}^{3}\left\|\mathbf{v}_{t}\right\|^{2}\right]+(2+8C^{2}+24C^{2}L^{2})\sigma^{2}.$
@@
$\leq$ $26C^{2}L^{2}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\mathbf{u}_{t}-g(\mathbf{x}_{t})\right\|^{2}\right]+\frac{8C^{2}L^{2}+24C^{4}L^{2}}{\beta}\mathbb{E}\left[\sum_{t=1}^{T}\eta_{t}^{2}\left\|\mathbf{v}_{t}\right\|^{2}\right]$
$\qquad+(2+8C^{2}+24C^{2}L^{2})\sigma^{2}T^{1/3}$
@@ Appendix D Proof of Theorem 4
$\begin{split}&\quad\ \mathbb{E}_{i_{t+1}}\left[\left\|\nabla F(\mathbf{x}_{t+1})-\mathbf{v}_{t+1}\right\|^{2}\right]\\ &=\mathbb{E}_{i_{t+1}}\left[\left\|(1-\beta)\mathbf{v}_{t}+\beta\mathbf{z}_{t+1}+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right)-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &=\mathbb{E}_{i_{t+1}}\left[\|(1-\beta)(\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t}))+\beta(\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1}))\right.\\ &\quad\left.+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right)\|^{2}\right]\\ &\leq\left\|(1-\beta)(\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t}))\right\|^{2}+\mathbb{E}_{i_{t+1}}\left[\left\|\beta(\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1}))\right.\right.\\ &\left.\left.\quad+(1-\beta)\left(\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right)\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\quad+2(1-\beta)^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})+\nabla F(\mathbf{x}_{t})-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]\\ &\quad+2(1-\beta)^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t+1})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right\|^{2}\right]\\ &\leq(1-\beta)^{2}\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]+2L^{2}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|^{2}\\ &\leq(1-\beta)\left\|\mathbf{v}_{t}-\nabla F(\mathbf{x}_{t})\right\|^{2}+2\beta^{2}\mathbb{E}_{i_{t+1}}\left[\left\|\mathbf{z}_{t+1}-\nabla F(\mathbf{x}_{t+1})\right\|^{2}\right]+2L^{2}\eta_{t}^{2}\left\|\mathbf{v}_{t}\right\|^{2}.\end{split}$ (18)
@@
$\leq$ $\mathbb{E}_{i_{t+1}}\left[(1+2n)L^{2}\eta_{t}^{3}\left\|\mathbf{v}_{t}\right\|^{2}\right.$
$\quad\left.+\eta_{t}(1+\frac{1}{2n})\left((1-\frac{1}{n})\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t})-g_{t}^{i_{t+1}}\right\|^{2}+\frac{1}{n}\left\|\nabla f_{i_{t+1}}(\mathbf{x}_{t})-\nabla f_{i_{t+1}}(\mathbf{x}_{t})\right\|^{2}\right)\right]$
@@
$\sum_{t=1}^{T}\eta_{t}^{2}L\left\|\mathbf{v}_{t}\right\|^{2}=$ ${L}\sum_{t=1}^{T}\frac{\left\|\mathbf{v}_{t}\right\|^{2}}{n^{1-\alpha}\left(\sum_{i=1}^{t}\left\|\mathbf{v}_{i}\right\|^{2}\right)^{2\alpha}}$
@@
$C_{3}=$ $2\Delta_{F}+\frac{14\alpha}{1-\alpha}\left(\frac{2L^{2}}{1-3\alpha}\left(\frac{24-72\alpha}{1-\alpha}\right)^{\frac{1-3\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{2\alpha}}$
$+\frac{\alpha}{1-\alpha}\left(\frac{L}{1-2\alpha}\left(\frac{4-8\alpha}{1-\alpha}\right)^{\frac{1-2\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{\alpha}}.$
@@
$\mathbb{E}\left[\sum_{t=1}^{T}\left\|\nabla F(\mathbf{x}_{t})-\mathbf{v}_{t}\right\|^{2}\right]$
$\leq$ $\frac{\alpha\sqrt{n}}{1-\alpha}\left(\frac{14L^{2}}{1-2\alpha}\left(\frac{2-4\alpha}{(1-\alpha)L}\right)^{\frac{1-2\alpha}{1-\alpha}}\right)^{\frac{1-\alpha}{\alpha}}+\frac{n^{\frac{\alpha}{2}}L}{2}\mathbb{E}\left[\left(\sum_{t=1}^{T}\left\|\mathbf{v}_{t}\right\|^{2}\right)^{1-\alpha}\right].$